Question 2 In the Keynesian cross, assume that the consumption function is given by C = 150 +0.7 (Y-T) Planned investment is:please answer part h to part k
Answer #1

1 h) Equation of planned expenditure: AE = C+It be = 150+0.(Y - 25) + 100-10(5) + 50 = 250 +0.77-175 = 2385 +0.7Y, At equilib

Interest rate -10 Output 800 500 009 700